OpenStudy (anonymous):

Hexagon A is a regular hexagon. The total length of all the sides of the hexagon is 24 inches. Hexagon A is dilated about its center to create Hexagon B. The length of each side of Hexagon B is inches. By what factor was Hexagon A dilated to create Hexagon B? I think it would be C. A. 4/5 B. 5/6 C. 1 1/5 D. 2

OpenStudy (anonymous):

Am I correct?

OpenStudy (mertsj):

Well. what is the length of the side of Hexagon B?

OpenStudy (anonymous):

4 4/5

OpenStudy (mertsj):

4x=4 4/5

OpenStudy (mertsj):

C

OpenStudy (mertsj):

Yes. You are correct. I was thinking octagon at first then I reread the problem

OpenStudy (anonymous):

\[\frac{4+(4/5)}{4}=\frac{6}{5}=1\frac{1}{5} \]C.

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

Isn't area of a hexagon = a^2 * ( 3 * sqrt(3) ) / 2, where a is the length of the hexagon. Since the length of the side increased from 4 to 4.8 so the area should increase by 1.44

OpenStudy (mertsj):

yw

OpenStudy (anonymous):

1.44 = square of (1 1/5)

OpenStudy (anonymous):

The question does not refer to areas or area ratios. The second hexagon can be constructed knowing how much larger the original side length is. The perimeter is proportional to the increase in side length

OpenStudy (anonymous):

Ok, I see, to me dilated meant area, but I see your point. @robtobey